Mg + 2 HCl = Mg Cl2 + H2
If 200 mL of Hydrogen were produced from 0.25g of Mg reacting, what is the percent yield
1
Expert's answer
2020-05-21T13:52:46-0400

Number of moles of H2 is equal to that of Mg in the reaction.

The amount of Mg reacted was

0.25 g / 24.3 g/mole = 0.0103 mole.

Amount of H2 produced at STP is

0.200 L / 22.4 L/mole = 0.0089 mole.

Percent yield is

0.0089/0.0103*100% = 86.7 %